# VR-GS: A Physical Dynamics-Aware Interactive Gaussian Splatting System in Virtual Reality

### v1

@@ 4.1. Gaussian Splatting
Gaussian splatting, proposed by Kerbl et al. (2023), is an explicit 3D representation to encapsulate 3D scene information using a set of 3D anisotropic Gaussian kernels, each with learnable mean $\mu$, opacity $\sigma$, covariance matrix $\Sigma$, and spherical harmonic coefficients $\mathcal{C}$. Spherical harmonics (SH) are expansions of the view-dependent color function defined on the unit sphere.
To render a view, the 3D splats are projected to 2D screen space ordered by their $z$-depth. The color $C$ of a pixel is computed by $\alpha$-blending of these 2D Gaussians from near to far:
@@ 4.2. VR-GS Assets Preparation
Segmentation
The segmented GS is constructed during the GS reconstruction. We first generate 2D masks on the multi-view RGB images by utilizing a 2D segmentation model (Cheng et al., 2023). Each segmented part is assigned a different color that is consistent across different views. Subsequently, we enhance the scene representation by integrating three additional learnable RGB attributes into the 3D Gaussian kernels. During the reconstruction process, each 3D Gaussian kernel will automatically learn what object it belongs to utilizing a segmentation loss function $L_{\text{seg}}$:
@@
Inpainting
After 3D GS segmentation, we extract all objects separately from the scene. This process of object removal, however, results in the emergence of holes within the regions that are previously occluded. To alleviate the issue, we utilize 2D inpainting tool LaMa (Suvorov et al., 2022) to guide the 3D inpainting of Gaussian kernels. We freeze the Gaussian kernels located outside of the holes and then use an inpainting loss $L_{\text{inpaint}}=L_{1}(I_{\text{inpainted}},I)$ to optimize a Gaussian kernel patch under the guidance of the 2D inpainted images $I_{\text{inpainted}}$ for the current 3D GS rendering $I$. The result of our inpainting is validated in Section 5.
Mesh Generation
Due to our design choice to use mesh-based simulation, we generate a tetrahedral mesh for each group for the segmented 3D Gaussian kernels. These meshes will not be rendered during the interaction but only serve as the media of dynamics. Hence, we do not require high-fidelity meshes, which, in return, boost the performance of our system. To construct a mesh, we first use internal filling proposed by (Xie et al., 2023) to fill particles into the void internal region, which failed to be reconstructed by GS. Then we treat centers of Gaussians as a point cloud and convert it to a voxel representation represented by the VDB format(Museth, 2013). A water-tight surface mesh is extracted from this VDB using marching cubes (Lorensen and Cline, 1987), which is then served as the boundary of the volume to generate a simulation-ready tetrahedral mesh by TetGen (Si, 2015).
@@ 4.3. Unified Framework for Simulation and Rendering
As derived in Xie et al. (2023), Gaussian kernels can be deformed by the deformation map given by the simulation. We follow the Gaussian kinematics of this work but replace the simulation with eXtended position-based dynamics (XPBD) (Macklin et al., 2016) to achieve real-time interactions. We employ the strain energy constraint as the elastic model and adopt the velocity-based damping model in the XPBD framework. To resolve collisions between meshes in real-time, penalty forces based on point-triangle distance constraint are added on close triangles pairs detected by LBVH (Pantaleoni and Luebke, 2010).
In mesh-based simulation, the deformation map is given by a piece-wise linear map, where the deformation gradient is hence constant within each tetrahedron. Given a tetrahedron with the rest-shape configuration $\{\bm{x}_{0}^{0},\bm{x}_{1}^{0},\bm{x}_{2}^{0},\bm{x}_{3}^{0}\}$ and the current configuration $\{\bm{x}_{0},\bm{x}_{1},\bm{x}_{2},\bm{x}_{3}\}$, the deformation gradient is defined as
@@
where $w_{i}$ is the barycentric coordinates of the initial center $\bm{\mu}_{0}$ in the rest-shape configuration, and $\bm{\Sigma}_{0}$ is the initial covariance matrix (Xie et al., 2023). The deformed Gaussian kernels can be directly rendered by the point splatting procedure. However, the direct embedding of Gaussian centers cannot guarantee that every ellipse shape is completely enveloped by some tetrahedron inside the simulation mesh. As shown in Figure 3, this can lead to spiky artifacts.
Observe that kernels that are completely inside the tetrahedron will always be inside. This motivates us to propose a two-level embedding procedure:
@@
Shadow Map
While the original global illumination of the scene can be accurately learned and baked by the spherical harmonics on each Gaussian, the shadow will no longer be aligned with the object when it is moving or deforming. The introduction of the shadow map (Wanger et al., 1992) into the GS framework can enhance the immersive experience in VR. More importantly, it can guide human perception of the spatial relationships between objects: during manipulation, users will rely on the shadow to determine the distance between objects. The shadow map is a fast real-time algorithm and is well-aligned with the GS rasterization pipeline. We follow Equation 1 to estimate the depth map from the light source and test the visibility for each Gaussian using this depth map. The influence of shadow maps is ablated in Section 5.
Figure 4. Inpainting Evaluation. GS struggles to reconstruct occluded surfaces. By leveraging LAMA (Suvorov et al., 2022), we produce 2D inpainted results that guide the 3D scene inpainting, enhancing the realism of the 3D representation.
@@ 5. Evaluation
The two-level embedding is a crucial component in our physical dynamics-aware system, integrating the tetrahedra cage mesh with the embedded Gaussian. In the inset figure, we conduct an ablation study to validate the effectiveness of our two-level embedding approach. Under conditions of extreme stretching or twisting, the naive embedding method, which merely embeds the Gaussian kernel within the closest tetrahedron, leads to severe spiky artifacts. Our two-level embedding strategy addresses this by initially embedding each Gaussian into a localized, independent tetrahedron, followed by embedding the vertices of this tetrahedron into the cage mesh. The deformation of each Gaussian kernel is determined by averaging the deformations at these vertices, resulting in a smoother deformation gradient field than the naive approach and significantly reducing spiky artifacts.
@@
With the guidance of 2D segmentation image results, our system achieves object-level 3D scene reconstruction, facilitating convenient post-physics-based manipulation. However, GS is limited to reconstructing surfaces visible in the provided multi-view training images. Consequently, some object and background areas, unseen in these images, remain unreconstructed by GS, leading to ’black hole’ like artifacts when foreground objects are moved. To address this, our system integrates the object segmentation mask with LAMA (Suvorov et al., 2022), a 2D inpainting model, to generate inpainted multi-view images. These images then guide the fine-tuning and inpainting of our 3D GS scene, yielding a more complete and realistic user experience, as evidenced in Figure 4.
@@
Furthermore, we take advantage of the shadow map to add extra time-dependent shadow into the GS scene, as shown in 5. Original GS models shadows as fixed surface textures, and thus fails to provide dynamic shadows as objects move around the scene. VR-GS allows users to choose the parameters, e.g. position and direction, of the light source to reproduce the lighting setting of the original scene, leading to a more realistic VR environment.
@@ 6. Experiment
In this section, we benchmark our VR-GS system’s simulation performance against other leading methods in GS and NeRF manipulation. Additionally, we present immersive and interactive demos within our VR-GS system. Our prototype, developed on Unity with a CUDA-implemented plugin, is tested using a Quest Pro Head-Mounted Display (HMD) and corresponding controllers, on a high-performance Intel(R) Core(TM) i9-14900KF CPU with 32GB memory and an NVIDIA GeForce RTX 4090 GPU.
@@ 6.1. Performance
We compared VR-GS against two state-of-the-art NeRF/GS physics-based manipulation methods: PAC-NeRF (Li et al., 2023a) and PhysGaussian (Xie et al., 2023). PAC-NeRF primarily concentrates on estimating material parameters from multi-view videos in reconstructed NeRF scenes. Although it offers novel dynamic generation capabilities, the resulting visuals often fall short in rendering quality. In contrast, PhysGaussian, leveraging GS, produces superior photorealistic dynamics. However, the Material Point Method (MPM) used by both systems can limit real-time performance in extensive scenes. For a fair comparison, we standardize the frame time ($\Delta t_{\text{frame}}=1/25$ sec) and the simulation step time ($\Delta t_{\text{step}}=0.0001$ sec) across all methods. For VR-GS, we set the XPBD iteration per substep to $1$, as the small $\Delta t_{\text{step}}$ sufficiently resolves the dynamics. As depicted in Figure 6 and the inset table, VR-GS not only matches the visual quality of PhysGaussian but also outperforms PAC-NeRF in clarity and realism. Crucially, our XPBD-based simulation framework allows for significantly faster simulations, making VR-GS ideal for real-time, physics-aware interactions in VR.
Figure 6. Visual Quality Comparison. Our method synthesizes competitive visual results compared to PhysGuassian (Xie et al., 2023) and significantly outperforms PAC-NeRF (Li et al., 2023a).
@@ 6.2. VR Demos
Figure 7. Fox, Bear, and Horse. Our proposed VR-GS allows users to manipulate 3D GS at a real-time rate with physically plausible response.
Figure 8. Timing Breakdown of Demos in Section 6.2.
Our VR-GS system’s capability to replicate real-world scenarios is showcased through several representative demos. These demos, employing GS and our dynamic simulation model, illustrate the system’s interactive and immersive qualities. Detailed simulation setups and a timing breakdown for these demos are provided in Table 1 and Figure 8.
@@
VR-GS empowers users to edit 3D Gaussian kernels in a user-friendly manner, utilizing our XPBD-based physics engine for dynamic manipulation. We demonstrate this through three examples: a fox, a bear, and a horse, reconstructed from the Instant-NGP dataset (Müller et al., 2022), the Instruct-NeRF2NeRF (Haque et al., 2023), and the Tanks and Temples dataset (Knapitsch et al., 2017), respectively. Each object’s Gaussian kernels are segmented from their scenes and converted into tetrahedron meshes, allowing for physics-based post-manipulation. The interactions, as depicted in Figure 7, exhibit lively responses to user inputs within the 3D environment.
Ring Toss
In this scenario, we demonstrate the system’s ability to seamlessly integrate new virtual objects into pre-existing scenes. A room scene, reconstructed from our collected real-world data, and a virtual ring, modeled in Blender, are used to create a ring toss game. The dining table and objects on it serve as collision boundaries for the ring. This interaction, shown in Figure 9, highlights VR-GS’s capacity to blend reconstructed and virtual objects in a unified framework, creating an immersive VR experience.
@@
( Figure 7) Fox 304,875 28,205 1 10 161.2
( Figure 7) Bear 2,525,891 60,590 10 10 36.3
( Figure 7) Horse 1,295,223 10,611 5 10 115.4
( Figure 9) Ring Toss 956,116 17,513 20 10 39.0
( Figure 10) Toy Collection 2,555,482 46,810 5 5 19.3
( Figure 4) Box Moving 1,769,412 1,434 10 1 73.8
(Figure 1) Animal Crossing 1,031,191 44,969 15 10 33.5
    #Iter: XPBD iteration count; CD: collision detection per step.
Figure 9. Ring Toss. Utilizing our VR-GS system, participants can engage in an interactive ring toss game simulated within authentic environments.
Toy Collection
To further assess the system’s efficiency and efficacy with multiple interactive objects, we reconstructed a living room scene and a set of plush toys, the latter partly synthesized using LucidDreamer (Liang et al., 2023). This setup, illustrated in Figure 10, allows users to realistically interact with the toys, showcasing VR-GS’s ability to handle complex scenarios involving a mix of 3D GS objects derived from real and synthesized environments.
Animal Crossing
The Animal Crossing background scene is the same as the Ring Toss scene. All four animal models sitting on chairs are reconstructed from Toy Collection. VR-GS permitted free deformation of the animal models, enabling users to freely interact with them. As demonstrated in Figure 1, this setup leveraged our sophisticated generative dynamics, resulting in objects reacting to impacts in a manner that closely mimics real-world physics.
Figure 10. Toy Collection. We place all the toys on the sofa, and then move them into the basket. After shaking the basket, all the toys fly out in a realistic way.

### v2

@@ 4.1. Gaussian Splatting
Gaussian splatting, proposed by Kerbl et al. [2023], is an explicit 3D representation to encapsulate 3D scene information using a set of 3D anisotropic Gaussian kernels, each with learnable mean $\mu$, opacity $\sigma$, covariance matrix $\Sigma$, and spherical harmonic coefficients $\mathcal{C}$. Spherical harmonics (SH) are expansions of the view-dependent color function defined on the unit sphere. To render a view, the 3D splats are projected to 2D screen space ordered by their $z$-depth. The color $C$ of a pixel is computed by $\alpha$-blending of these 2D Gaussians from near to far:
@@ 4.2. VR-GS Assets Preparation
4.2.1. Segmentation
The segmented GS is constructed during GS training. We first generate 2D masks on the multi-view RGB images by utilizing a 2D segmentation model [Cheng et al., 2023]. Each segmented part is assigned a different color that is consistent across different views. Subsequently, we enhance the scene representation by integrating three additional learnable RGB attributes into the 3D Gaussian kernels. During the reconstruction process, each 3D Gaussian kernel will automatically learn what object it belongs to utilizing a segmentation loss function $L_{\text{seg}}$:
@@ 4.2.1. Segmentation
4.2.2. Inpainting
After 3D GS segmentation, we extract all objects separately from the scene. This process of object removal, however, results in the emergence of holes within the regions that are previously occluded. To alleviate the issue, we utilize a 2D inpainting tool LaMa [Suvorov et al., 2022] to guide the 3D inpainting of Gaussian kernels. We freeze the Gaussian kernels located outside of the holes and then use an inpainting loss $L_{\text{inpaint}}=L_{1}(I_{\text{inpainted}},I)$ to optimize a Gaussian kernel patch under the guidance of the 2D inpainted images $I_{\text{inpainted}}$ for the current 3D GS rendering $I$. The result of our inpainting is validated in Section 5.
4.2.3. Mesh Generation
Due to our design choice to use mesh-based simulation, we generate a tetrahedral mesh for each group of the segmented 3D GS kernels. These meshes will not be rendered during the interaction but only serve as the media of dynamics. Hence we can use a moderate mesh resolution which does not hinder performance. To construct a simulation mesh, we first use internal filling proposed by [Xie et al., 2024] to fill particles into the void internal region that is not reconstructed by GS. Then we treat centers of Gaussians as a point cloud and convert it to a voxelized VDB representation [Museth, 2013]. A water-tight surface mesh is then extracted using marching cubes [Lorensen and Cline, 1987] and tetrahedralized into a finite element mesh using TetGen [Si, 2015].
@@ 4.3. Unified Framework for Simulation and Rendering
As shown in Xie et al. [2024], Gaussian kernels can be deformed by the simulated deformation field. We follow the Gaussian kinematics of this work but replace the simulator with XPBD [Macklin et al., 2016] to achieve real-time interactions. We employ the strain energy constraint as the elastic model and adopt the velocity-based damping model in the XPBD framework.
4.3.1. Physical Parameters
The physical parameters, such as densities and material stiffness, are tunable hyperparameters. Following PhysGaussian [Xie et al., 2024], we manually set physical parameters for all interactable objects, including Young’s modules ($E$), Poisson ratio ($\nu$), and density ($\rho$). We start from an initial configuration $E=1000\text{ Pa},\nu=0.3,\rho=1000\text{ kg}$ for each object, then tune them to produce visually plausible dynamics. The Young’s modulus governs the overall stiffness of an object and is manually adjusted by analyzing the severity of deformation of its free fall collision with the ground. Additionally, we adjust the relative density between two objects based on their deformations under collision. In practice we usually arrive at a finalized parameter setting that gives visually plausible results within 10 iterations of this manual process.
4.3.2. Embedding
In our mesh-based simulation, the deformation map is piece-wise linear, with the resulting deformation gradient piece-wise constant within each tetrahedron. Given a tetrahedron with the rest-shape configuration $\{\bm{x}_{0}^{0},\bm{x}_{1}^{0},\bm{x}_{2}^{0},\bm{x}_{3}^{0}\}$ and the current configuration $\{\bm{x}_{0},\bm{x}_{1},\bm{x}_{2},\bm{x}_{3}\}$, the deformation gradient is defined as
@@ 4.3.2. Embedding
where $w_{i}$ is the barycentric coordinates of the initial center $\bm{\mu}_{0}$ in the rest-shape configuration, and $\bm{\Sigma}_{0}$ is the initial covariance matrix [Xie et al., 2024]. The deformed Gaussian kernels can be directly rendered by the point splatting procedure. However, the direct embedding of Gaussian centers cannot guarantee that every ellipse shape is completely enveloped by some tetrahedron inside the simulation mesh. As shown in Figure 3, this can lead to spiky artifacts. Observe that kernels that are completely inside the tetrahedron will always be inside. This motivates us to propose a two-level embedding procedure:
  1. (1)
@@
4.3.3. Shadow Map
While the original global illumination of the scene can be accurately learned and baked by the spherical harmonics on each Gaussian, the shadow will no longer be aligned with the object when it is moving or deforming. Bringing shadow map [Wanger et al., 1992] into the GS framework can enhance the immersive experience in VR. More importantly, it can guide human perception of the spatial relationships between objects: during manipulation, users will rely on the shadow to determine the distance between objects. The shadow map is a fast real-time algorithm and is well-aligned with the GS rasterization pipeline. We follow Equation 1 to estimate the depth map from the light source and test the visibility for each Gaussian using this depth map. The influence of shadow maps is studied in Section 5.
Figure 4. Two-level Embedding Evaluation. Our two-level embedding alleviates the commonly seen spiky artifacts for deformed GS kernels.
Figure 5. Inpainting Evaluation. GS struggles to reconstruct occluded surfaces. By leveraging LAMA [Suvorov et al., 2022], we produce 2D inpainted results that guide the 3D scene inpainting, enhancing the realism of the 3D representation.
@@ 5. Evaluation
The two-level embedding is a crucial component in our physical dynamics-aware system, integrating the tetrahedra cage mesh with the embedded Gaussian. In Figure 4, we conduct an ablation study to validate the effectiveness of our two-level embedding approach. Under conditions of extreme stretching or twisting, the naive embedding method, which simply embeds the Gaussian kernel within the closest tetrahedron, leads to severe spiky artifacts. Our two-level embedding strategy addresses this by initially embedding each Gaussian into a localized, independent tetrahedron, followed by embedding the vertices of this tetrahedron into the cage mesh. The deformation of each Gaussian kernel is determined by averaging the deformations at these vertices, resulting in a smoother deformation gradient field than the naive approach and significantly reducing spiky artifacts.
@@
With the guidance of 2D image segmentation results, our system achieves object-level 3D scene reconstruction, facilitating convenient post-physics-based manipulation. However, GS is limited to reconstructing surfaces visible in the provided multi-view training images. Consequently, some object and background areas, unseen in these images, remain unreconstructed by GS, leading to “black hole” like artifacts when foreground objects are moved. To address this, our system integrates the object segmentation mask with LAMA [Suvorov et al., 2022], a 2D inpainting model, to generate inpainted multi-view images. These images then guide the fine-tuning and inpainting of our 3D GS scene, yielding a more complete and realistic user experience, as evidenced in Figure 5.
@@
Furthermore, we take advantage of the shadow map to add extra time-dependent shadows into the GS scene, as shown in 6. Original GS represents shadows as textures and thus fails to provide dynamic shadows when objects move. VR-GS allows users to choose the parameters, e.g. position and direction, of the light source to reproduce the lighting setting of the original scene, leading to a more realistic VR environment.
@@ 6. Experiment
In this section, we benchmark our VR-GS system’s simulation performance against other methods in GS and NeRF manipulation. Additionally, we showcase interactive demonstrations on VR devices. Our prototype, developed in Unity with a CUDA-implemented plugin, is tested using a Quest Pro Head-Mounted Display (HMD) and corresponding controllers, on an Intel Core i9-14900KF CPU with 32GB memory and an NVIDIA GeForce RTX 4090 GPU.
@@ 6.1. Performance
VR-GS is designed for real-time, physics-aware interactions by integrating Gaussian kernels within a cage mesh for simulation. This cage mesh is derived from the VDB representation of Gaussians, with adjustable mesh resolution to influence quality. We first conduct experiments to explore the trade-offs between mesh quality and system performance. As shown in Figure 12, using a coarse cage mesh facilitates higher frame rates but may lead to the loss of fine details. Conversely, a higher-resolution mesh inevitably increases the computational cost. Moreover, XPBD also requires much more iterations to achieve convergence for finer meshes. Without sufficient iterations, the simulated object may appear to be overly soft, yielding unrealistic dynamics. In practice, we constrain the number of mesh vertices to between 10K and 30K to maintain a balance between high frame rates and accurate physical dynamics.
@@
We then compare VR-GS against two state-of-the-art NeRF/GS physics-based manipulation methods: PAC-NeRF [Li et al., 2023a] and PhysGaussian [Xie et al., 2024]. PAC-NeRF primarily concentrates on estimating material parameters from multi-view videos in reconstructed NeRF scenes. Although it offers novel dynamic generation capabilities, the resulting visuals often fall short in rendering quality. In contrast, PhysGaussian, leveraging GS, produces superior photorealistic dynamics. However, the Material Point Method (MPM) used by both systems can limit real-time performance in complex scenes. For a fair comparison, we standardize the frame time ($\Delta t_{\text{frame}}=1/25$ sec) and the simulation step time ($\Delta t_{\text{step}}=0.0001$ sec) across all methods. For VR-GS, we set the XPBD iteration per substep to $1$, as the small $\Delta t_{\text{step}}$ sufficiently resolves the dynamics. As depicted in Figure 7 and the inset table, VR-GS not only matches the visual quality of PhysGaussian but also outperforms PAC-NeRF in clarity and realism. Crucially, our XPBD-based simulation framework allows for significantly faster simulations, making VR-GS ideal for real-time physics-aware interactions.
Figure 7. Visual Quality Comparison. Our method synthesizes competitive visual results compared to PhysGuassian [Xie et al., 2024] and significantly outperforms PAC-NeRF [Li et al., 2023a].
@@ 6.2. VR Demos
We then showcase VR-GS’s capability to replicate real-world scenarios through several representative demos in VR. Detailed simulation setups and a timing breakdown for these demos are provided in Table 1 and Figure 13.
@@
VR-GS empowers users to edit 3D Gaussian kernels intuitively and efficiently, utilizing our XPBD-based physics engine for dynamic manipulation. We demonstrate this through three examples as depicted in Figure 9: a fox, a bear, and a horse, reconstructed from the Instant-NGP dataset [Müller et al., 2022], the Instruct-NeRF2NeRF [Haque et al., 2023], and the Tanks and Temples dataset [Knapitsch et al., 2017], respectively.
Ring Toss and Table Brick Game
In this example, we demonstrate the system’s ability to seamlessly integrate new virtual objects into existing scenes. We use a room scene reconstructed from real-world footage and virtual objects (rings and bricks modeled in Blender) to create ring tossing and table brick game (Figure 10). The dining table and objects on it serve as collision boundaries for the elastic ring and rigid bricks [Deul et al., 2016].
@@
(Figure 9) Fox 304,875 28,205 1 / 161.2
(Figure 9) Bear 2,525,891 19,472 5 / 76.9
(Figure 9) Horse 1,295,223 10,611 5 / 115.4
(Figure 10) Ring Toss 1,456,209 9,002 10 10 37.7
(Figure 10) Table Brick Game 1,866,835 33,279 10 10 41.4
(Figure 11) Toy Collection 1,665,128 46,428 20 20 24.3
(Figure 5) Box Moving 1,769,412 1,434 10 / 73.8
(Figure 1) Animal Crossing 1,312,670 36,386 10 5 34.7
(Figure 8) Just Dance 1,059,054 13,936 50 / 33.9
*#Iter: XPBD iterations per substep; CD: collision detections per step.
Toy Collection and Animal Crossing
We reconstructed a living room scene and a set of plush toys (Figure 11) for users to interact with. The yellow spherical toy is generated using a text-to-3D generator LucidDreamer [Liang et al., 2023], while the others are reconstructed. Additionally, four animal plush toys were placed on chairs, allowing users to manipulate and deform them freely (Figure 1).
Just Dance
VR-GS also supports animation of reconstructed GS humans (Figure 8). Utilizing a reconstructed human character, we leverage Mixamo [^1] https://www.mixamo.com/’s auto-rig to generate motion sequences of its surface mesh. These sequences then serve as the boundary condition for the simulation, finally producing the GS human animation that can be viewed in our immersive system.
Figure 8. Just Dance. VR-GS generates high-fidelity dance given a reconstructed human body.
6.3. Modeling Details
We employ the Toy Collection (Figure 11) as a case study to demonstrate the time and labor required to create a fully interactive VR environment from scratch. The setup process is as follows:
  1. (1)
    The initial stage involves setting up the real-world scene. For the Toy Collection, the indoor scene was captured directly as one scene, with all toys separately hung with strings as another scene. We completed it within 20 minutes.
  2. (2)
    Next, we capture multi-view images for Gaussian Splatting by recording videos of the two scenes with a consumer-level camera and converting them into image sequences. Camera intrinsic and extrinsic parameters are collected using COLMAP, with the entire process taking around 60 minutes, of which COLMAP occupies approximately 40 minutes.
  3. (3)
    We then proceed to image processing for segmentation and inpainting. Object classes are manually designated in the first video frame. Subsequent frames are automatically segmented using the method from Cheng et al. [2023], taking about 2 minutes. The inpainting of what’s below the basket is fully automated [Suvorov et al., 2022] and requires 5 minutes.
  4. (4)
    GS reconstruction is performed on the processed images, undergoing 30,000 training steps as specified by Kerbl et al. [2023] without additional treatment for the number of Gaussians, which takes about 30 minutes on a single RTX 4090.
  5. (5)
    After generating 3D GS, we execute internal filling, VDB reconstruction, and TetGen, all within 10 minutes.
  6. (6)
    Lastly, object world positions and physical parameters are specified and estimated in Unity, as detailed in Section 4.3.1, taking 20 minutes.
In summary, the total preparation time for the Toy Collection is approximately 2 hours and 30 minutes. Manual tasks include scene staging, video recording, object class specification, and physical parameter selection. All other processes are automated. Once modeled, the scene is ready for a physics-aware interactive VR experience.